# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ Acknowledgements.
Special thanks as well to Victoria Lee for proofreading.
@@ Appendix A Derivation of beam tracing equations
$\mathcal{P}\alpha^{(0)}\Big((\delta_{il}\delta_{jm}-\delta_{ij}\delta_{ml})\partial_{m}\phi^{(0)}\partial_{l}\phi^{(0)}+\epsilon_{ij}\Big)\hat{e}_{j}=0$ (102)
@@
By contracting with the conjugate polarisation, $\hat{e}^{*}_{i}$, we can eliminate the term involving the first order correction to the amplitude, as it is contracted by $\hat{e}^{*}_{i}D_{ij}=0$. We will define $(\bm{\nabla}_{\bm{K}})_{i}=\partial K_{i}$. Utilising the fact that:
@@
$\partial K_{\mu}D_{im}=\delta_{i\mu}K_{m}+\delta_{\mu m}K_{i}-2\delta_{im}K_{\mu}=(\delta_{il}\delta_{jm}-\delta_{im}\delta_{jl})(\delta_{j\mu}\delta_{l\nu}+\delta_{j\nu}\delta_{l\mu})K_{\nu},$ (107)
along with 103, we can manipulate our equation into the following form:
$\Bigg(\partial K_{l}(H)\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)+w_{l}\partial_{l}(H)\Bigg)\alpha^{(0)}=0.$ (108)
@@
$w_{l}\left(\frac{1}{g}\frac{\mathrm{d}K_{l}}{\mathrm{d}\tau}\hat{g}_{m}\partial K_{m}H+\partial_{l}H\right)=0.$ (109)
@@
$g_{i}\partial_{i}H+\frac{\mathrm{d}K_{l}}{\mathrm{d}\tau}\partial K_{i}H=\frac{\mathrm{d}H}{\mathrm{d}\tau}=0.$ (112)
@@
$\begin{split}\partial K_{i}H&=\frac{\mathrm{d}q_{i}}{\mathrm{d}\tau},\\ \partial_{i}H&=-\frac{\mathrm{d}K_{i}}{\mathrm{d}\tau}.\end{split}$ (113)
@@
$w_{k}\partial_{k}D_{ij}\hat{e}_{j}\mathcal{P}\alpha^{(0)}+\mathcal{P}\alpha^{(0)}\partial K_{l}D_{im}\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\hat{e}_{m}=-D_{ij}A_{j}^{(1)}.$ (114)
@@
$w_{k}(\partial_{k}H\hat{e}_{i}-D_{ij}\partial_{k}\hat{e}_{j})\mathcal{P}\alpha^{(0)}+\mathcal{P}\alpha^{(0)}(\partial K_{l}H\hat{e}_{i}-D_{ij}\partial K_{l}\hat{e}_{j})\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)=-D_{ij}A_{j}^{(1)}.$ (116)
@@
$w_{k}(-\frac{\mathrm{d}K_{k}}{\mathrm{d}\tau}\hat{e}_{i}-D_{ij}\partial_{k}\hat{e}_{j})\mathcal{P}\alpha^{(0)}+\mathcal{P}\alpha^{(0)}(g_{l}\hat{e}_{i}-D_{ij}\partial K_{l}\hat{e}_{j})\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)=-D_{ij}A_{j}^{(1)},$ (117)
@@
$D_{ij}\mathcal{P}\alpha^{(0)}\left(w_{k}\partial_{k}\hat{e}_{j}+\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\partial K_{l}\hat{e}_{j}\right)=D_{ij}A_{j}^{(1)}.$ (118)
@@
$\mathcal{P}\alpha^{(0)}\left(w_{k}\partial_{k}\hat{e}_{j}+\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\partial K_{l}\hat{e}_{j}\right)=A_{j}^{(1)}.$ (120)
@@
$\left(w_{k}\partial_{k}+\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\partial K_{l}\right)\hat{e}^{(0)}_{j}=e_{j}^{(1)}.$ (121)
@@
$(\alpha^{(1)}_{j})_{\mathrm{original}}=w_{k}(\Psi_{kl}\partial K_{l}+\partial_{k})\hat{e}_{j}\alpha^{(0)},$ (123)
which is akin to $\bm{A}^{(1)}$ in a previous paper on Gaussian beam tracingHallChen2022. Then, $(\partial_{i}A_{m})^{(1)}$ is given by:
$(\partial_{j}A_{m})^{(1)}=\partial_{j}(\mathcal{P})^{(1)}\alpha^{(0)}_{m}+\mathcal{P}(\partial_{j}(\alpha^{(0)}_{m}+({\alpha_{m}^{(1)}})_{\mathrm{original}}))^{(0)}-\mathrm{i}(\partial K_{l}\hat{e}_{m}\partial_{j}\partial_{l}(\mathcal{P})^{(0)})\alpha_{m}^{(0)},$ (124)
@@
$\partial_{\nu}(\alpha^{(0)}_{m}+({\alpha_{m}^{(1)}})_{\mathrm{original}}))^{(0)}=\frac{\hat{g}_{\nu}}{g}\left(\frac{\mathrm{d}\alpha^{(0)}}{\mathrm{d}\tau}\hat{e}_{m}+\alpha^{(0)}\frac{\mathrm{d}\hat{e}_{m}}{\mathrm{d}\tau}\right)+(\delta_{\nu k}-\hat{g}_{\nu}\hat{g}_{k})(\Psi_{kl}\partial K_{l}+\partial_{k})\hat{e}_{m}\alpha^{(0)}.$ (125)
This is due to the dependence of $({\alpha_{m}^{(1)}})_{\mathrm{original}}$ on $\bm{w}$, whereas $\alpha^{(0)}$ only depends on $\tau$. Substituting this into our second-order expansion of the Helmholtz equation, we have:
$\begin{split}\hat{e}^{*}_{i}\Bigg(&(\delta_{il}\delta_{jm}-\delta_{im}\delta_{jl})(\delta_{j\mu}\delta_{l\nu}+\delta_{j\nu}\delta_{l\mu})\\ &\Big(-K_{\mu}\Psi_{\nu k}w_{k}\big(\overbrace{-\mathrm{i}\partial K_{l}\hat{e}_{m}\partial_{l}(\mathcal{P})^{(0)}\alpha^{(0)}}^{\mathrm{1}}+\mathcal{P}(\alpha^{(1)}_{m})_{\mathrm{original}}\big)\\ &-(\frac{1}{2}\Psi_{\mu k}w_{k}\Psi_{\nu j}w_{j}+K_{\mu}\partial_{\nu}\phi^{(2)})\hat{e}_{m}\mathcal{P}\alpha^{(0)}\\ &\overbrace{+\mathrm{i}(\Psi_{\mu k}w_{k}\partial_{\nu}(\mathcal{P})^{(0)}\hat{e}_{m})\alpha^{(0)}}^{\mathrm{1}}\\ &+\mathrm{i}K_{\mu}\big(\mathcal{P}\partial_{\nu}(\alpha^{(0)}_{m}+(\alpha^{(1)}_{m})_{\mathrm{original}})+\overbrace{\partial_{\nu}\mathcal{P}^{(1)}\alpha^{(0)}_{m}}^{\mathrm{4}}+\overbrace{\partial_{\nu}\mathcal{P}^{(0)}(\alpha^{(1)}_{m})_{\mathrm{original}}}^{\mathrm{1+3}}\overbrace{-\mathrm{i}\partial K_{l}\hat{e}_{m}\partial_{l}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}}^{2}\big)\\ &+\mathrm{i}\frac{1}{2}\Psi_{\mu\nu}\hat{e}_{m}\mathcal{P}\alpha^{(0)}+\overbrace{\frac{1}{2}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\hat{e}_{m}\alpha^{(0)}}^{2}\Big)\Bigg)\\ &=\hat{e}^{*}_{i}w_{k}\partial_{k}D_{ij}\Big(\mathcal{P}(\alpha^{(1)}_{j})_{\mathrm{original}}\overbrace{-\mathrm{i}\partial K_{l}\hat{e}_{j}\partial_{l}(\mathcal{P})^{(0)}\alpha^{(0)}}^{3}\Big)+\frac{1}{2}\hat{e}^{*}_{i}w_{k}w_{l}\partial_{k}\partial_{l}D_{ij}\hat{e}_{j}\mathcal{P}\alpha^{(0)}\end{split}$ (126)
@@
$(1):\mathrm{i}\hat{e}^{*}_{i}\big(\partial K_{\mu}\partial K_{\nu}(D_{ij}\hat{e}_{j})\big)w_{k}\Psi_{k\mu}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}$ (128)
$(2):\hat{e}^{*}_{i}\big(\partial K_{\mu}\partial K_{\nu}(D_{ij}\hat{e}_{j})\big)\frac{1}{2}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}$ (129)
$(3):\mathrm{i}\hat{e}^{*}_{i}\big(\partial_{\mu}\partial K_{\nu}(D_{ij}\hat{e}_{j})\big)w_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}$ (130)
@@
$\mathrm{i}\hat{e}^{*}_{i}(\partial K_{\nu}D_{im}\Psi_{\nu k}w_{k}\partial K_{l}\hat{e}_{m}\partial_{l}(\mathcal{P})^{(0)}+\partial K_{\mu}\partial K_{\nu}D_{im}\Psi_{\mu k}w_{k}\hat{e}_{m}\partial_{\nu}(\mathcal{P})^{(0)}+\partial K_{\nu}D_{im}\Psi_{lk}w_{k}\partial K_{l}\hat{e}_{m}\partial_{\nu}(\mathcal{P})^{(0)})\alpha^{(0)}.$ (131)
@@
$\frac{1}{2}\hat{e}^{*}_{i}(\partial K_{\mu}\partial K_{\nu}D_{im}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\hat{e}_{m}+2\partial K_{\nu}D_{im}\partial_{l}\partial_{\nu}(\mathcal{P})^{(0)}\partial K_{l}\hat{e}_{m})\alpha^{(0)}.$ (132)
@@
$\mathrm{i}\hat{e}^{*}_{i}(\partial K_{\nu}D_{im}w_{k}\partial_{k}\hat{e}_{m}\partial_{\nu}(\mathcal{P})^{(0)}+\partial_{k}D_{im}w_{k}\partial K_{l}\hat{e}_{m}\partial_{l}(\mathcal{P})^{(0)})\alpha^{(0)}.$ (133)
Since $\partial_{\mu}\partial K_{\nu}D_{ij}=0$, relabelling dummy indices gets us the expression in eq. 130. Thus, all the terms involving derivatives of the profile function add up to the following expression:
$\begin{split}\mathrm{i}&\Big(\partial_{\tau}(\mathcal{P})+w_{y}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\partial_{w_{x}}(\mathcal{P})+w_{x}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\partial_{w_{y}}(\mathcal{P})\Big)\alpha^{(0)}+\mathrm{i}\hat{e}^{*}_{i}\big(\partial K_{\mu}\partial K_{\nu}(D_{ij}\hat{e}_{j})\big)w_{k}\Psi_{k\mu}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}\\ &+\hat{e}^{*}_{i}\big(\partial K_{\mu}\partial K_{\nu}(D_{ij}\hat{e}_{j})\big)\frac{1}{2}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}+\mathrm{i}\hat{e}^{*}_{i}\big(\partial_{\mu}\partial K_{\nu}(D_{ij}\hat{e}_{j})\big)w_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\alpha^{(0)}.\end{split}$ (134)
@@
$\frac{\mathrm{d}}{\mathrm{d}\tau}(T_{\mu_{0}\mu_{1}\mu_{2}...\mu_{n}}):=\frac{\mathrm{d}T_{\mu_{0}\mu_{1}\mu_{2}...\mu_{n}}}{\mathrm{d}\tau}+\sum_{i=0}^{n}\sum_{j=0}^{2}\hat{x}_{j}^{\mu_{n}}\frac{\mathrm{d}\hat{x}_{j}^{\rho}}{\mathrm{d}\tau}T_{\mu_{0}...\mu_{n-1}\rho\mu_{n+1}...\mu_{n}}.$ (135)

### v3

@@ Acknowledgements.
J Ruiz Ruiz was funded by the Engineering and Physical Sciences Research Council (EPSRC), with grant numbers EP/R034737/1 and EP/W026341/1.
@@ Appendix A Derivation of beam tracing equations
In this appendix, we will utilise Einstein summation convention throughout. As per convention, the symbol $\partial_{\mu}$’ refers to the partial derivative with respect to a given coordinate chart $r_{\mu}$, which in this case is any chosen coordinate chart for our three-dimensional Euclidean space. $\frac{\partial}{\partial K_{\mu}}$’ refers to the partial derivative of a function with respect to $K_{\mu}$. This is to facilitate simpler manipulation of the numerous contractions and identities we will be introducing in the proceeding appendix. At the end, the results we obtain can easily be expressed in tensor notation.
@@
$\mathcal{P}A^{(0)}\Big((\delta_{il}\delta_{jm}-\delta_{ij}\delta_{ml})\partial_{m}\phi^{(0)}\partial_{l}\phi^{(0)}+\epsilon_{ij}\Big)\hat{e}_{j}=0$ (93)
@@
By contracting with the conjugate polarisation, $\hat{e}^{*}_{i}$, we can eliminate the term involving the first order correction to the amplitude, as it is contracted by $\hat{e}^{*}_{i}D_{ij}=0$. We will define $(\bm{\nabla}_{\bm{K}})_{i}=\frac{\partial}{\partial K_{i}}$. Utilising the fact that:
@@
$\frac{\partial}{\partial K_{\mu}}D_{im}=\delta_{i\mu}K_{m}+\delta_{\mu m}K_{i}-2\delta_{im}K_{\mu}=(\delta_{il}\delta_{jm}-\delta_{im}\delta_{jl})(\delta_{j\mu}\delta_{l\nu}+\delta_{j\nu}\delta_{l\mu})K_{\nu},$ (98)
along with 94, we can manipulate our equation into the following form:
$\Bigg(\frac{\partial}{\partial K_{l}}(H)\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)+w_{l}\partial_{l}(H)\Bigg)A^{(0)}=0.$ (99)
@@
$w_{l}\left(\frac{1}{g}\frac{\mathrm{d}K_{l}}{\mathrm{d}\tau}\hat{g}_{\mu}\frac{\partial}{\partial K_{\mu}}H+\partial_{l}H\right)=0.$ (100)
@@
$g_{i}\partial_{i}H+\frac{\mathrm{d}K_{l}}{\mathrm{d}\tau}\frac{\partial}{\partial K_{i}}H=\frac{\mathrm{d}H}{\mathrm{d}\tau}=0.$ (103)
@@
$\begin{split}\frac{\partial}{\partial K_{i}}H&=\frac{\mathrm{d}q_{i}}{\mathrm{d}\tau},\\ \partial_{i}H&=-\frac{\mathrm{d}K_{i}}{\mathrm{d}\tau}.\end{split}$ (104)
@@
$w_{k}\partial_{k}D_{ij}\hat{e}_{j}\mathcal{P}A^{(0)}+\mathcal{P}A^{(0)}\frac{\partial}{\partial K_{l}}D_{im}\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\hat{e}_{m}=-D_{ij}A_{j}^{(1)}.$ (105)
@@
$w_{k}(\partial_{k}H\hat{e}_{i}-D_{ij}\partial_{k}\hat{e}_{j})\mathcal{P}A^{(0)}+\mathcal{P}A^{(0)}\left(\frac{\partial}{\partial K_{l}}H\hat{e}_{i}-D_{ij}\frac{\partial}{\partial K_{l}}\hat{e}_{j}\right)\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)=-D_{ij}A_{j}^{(1)}.$ (107)
@@
$w_{k}\left(-\frac{\mathrm{d}K_{k}}{\mathrm{d}\tau}\hat{e}_{i}-D_{ij}\partial_{k}\hat{e}_{j}\right)\mathcal{P}A^{(0)}+\mathcal{P}A^{(0)}\left(g_{l}\hat{e}_{i}-D_{ij}\frac{\partial}{\partial K_{l}}\hat{e}_{j}\right)\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)=-D_{ij}A_{j}^{(1)},$ (108)
@@
$D_{ij}\mathcal{P}A^{(0)}\left(w_{k}\partial_{k}\hat{e}_{j}+\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\frac{\partial}{\partial K_{l}}\hat{e}_{j}\right)=D_{ij}A_{j}^{(1)}.$ (109)
@@
$\mathcal{P}A^{(0)}\left(w_{k}\partial_{k}\hat{e}_{j}+\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\frac{\partial}{\partial K_{l}}\hat{e}_{j}\right)=A_{j}^{(1)}.$ (111)
@@
$\left(w_{k}\partial_{k}+\Big(\Psi_{lk}w_{k}-\mathrm{i}(\partial_{l}(\ln\mathcal{P}))^{(0)}\Big)\frac{\partial}{\partial K_{l}}\right)\hat{e}^{(0)}_{j}=e_{j}^{(1)}.$ (112)
@@
$(\alpha^{(1)}_{j})_{\mathrm{original}}=w_{k}\left(\Psi_{kl}\frac{\partial}{\partial K_{l}}+\partial_{k}\right)\hat{e}_{j}A^{(0)},$ (114)
which is akin to $\bm{A}^{(1)}$ in a previous paper on Gaussian beam tracing[HallChen2022]. Then, $(\partial_{i}A_{m})^{(1)}$ is given by:
$(\partial_{j}A_{m})^{(1)}=\partial_{j}(\mathcal{P})^{(1)}A^{(0)}_{m}+\mathcal{P}(\partial_{j}(A^{(0)}_{m}+({\alpha_{m}^{(1)}})_{\mathrm{original}}))^{(0)}-\mathrm{i}\left(\frac{\partial}{\partial K_{l}}\hat{e}_{m}\partial_{j}\partial_{l}(\mathcal{P})^{(0)}\right)\alpha_{m}^{(0)},$ (115)
@@
$\partial_{\nu}(A^{(0)}_{m}+({\alpha_{m}^{(1)}})_{\mathrm{original}}))^{(0)}=\frac{\hat{g}_{\nu}}{g}\left(\frac{\mathrm{d}A^{(0)}}{\mathrm{d}\tau}\hat{e}_{m}+A^{(0)}\frac{\mathrm{d}\hat{e}_{m}}{\mathrm{d}\tau}\right)+\left(\delta_{\nu k}-\hat{g}_{\nu}\hat{g}_{k})(\Psi_{kl}\frac{\partial}{\partial K_{l}}+\partial_{k}\right)\hat{e}_{m}A^{(0)}.$ (116)
This is due to the dependence of $({\alpha_{m}^{(1)}})_{\mathrm{original}}$ on $\bm{w}$, whereas $A^{(0)}$ only depends on $\tau$. Substituting this into our second-order expansion of the Helmholtz equation, we have:
$\begin{split}\hat{e}^{*}_{i}\Bigg(&(\delta_{il}\delta_{jm}-\delta_{im}\delta_{jl})(\delta_{j\mu}\delta_{l\nu}+\delta_{j\nu}\delta_{l\mu})\\ &\Big(-K_{\mu}\Psi_{\nu k}w_{k}\Big(\overbrace{-\mathrm{i}\frac{\partial}{\partial K_{l}}\hat{e}_{m}\partial_{l}(\mathcal{P})^{(0)}A^{(0)}}^{\mathrm{1}}+\mathcal{P}(\alpha^{(1)}_{m})_{\mathrm{original}}\Big)\\ &-(\frac{1}{2}\Psi_{\mu k}w_{k}\Psi_{\nu j}w_{j}+K_{\mu}\partial_{\nu}\phi^{(2)})\hat{e}_{m}\mathcal{P}A^{(0)}\\ &\overbrace{+\mathrm{i}(\Psi_{\mu k}w_{k}\partial_{\nu}(\mathcal{P})^{(0)}\hat{e}_{m})A^{(0)}}^{\mathrm{1}}\\ &+\mathrm{i}K_{\mu}\Big(\mathcal{P}\partial_{\nu}(A^{(0)}_{m}+(\alpha^{(1)}_{m})_{\mathrm{original}})+\overbrace{\partial_{\nu}\mathcal{P}^{(1)}A^{(0)}_{m}}^{\mathrm{4}}+\overbrace{\partial_{\nu}\mathcal{P}^{(0)}(\alpha^{(1)}_{m})_{\mathrm{original}}}^{\mathrm{1+3}}\overbrace{-\mathrm{i}\frac{\partial}{\partial K_{l}}\hat{e}_{m}\partial_{l}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}}^{2}\Big)\\ &+\mathrm{i}\frac{1}{2}\Psi_{\mu\nu}\hat{e}_{m}\mathcal{P}A^{(0)}+\overbrace{\frac{1}{2}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\hat{e}_{m}A^{(0)}}^{2}\Big)\Bigg)\\ &=\hat{e}^{*}_{i}w_{k}\partial_{k}D_{ij}\Big(\mathcal{P}(\alpha^{(1)}_{j})_{\mathrm{original}}\overbrace{-\mathrm{i}\frac{\partial}{\partial K_{l}}\hat{e}_{j}\partial_{l}(\mathcal{P})^{(0)}A^{(0)}}^{3}\Big)+\frac{1}{2}\hat{e}^{*}_{i}w_{k}w_{l}\partial_{k}\partial_{l}D_{ij}\hat{e}_{j}\mathcal{P}A^{(0)}\end{split}$ (117)
@@
$(1):\mathrm{i}\hat{e}^{*}_{i}\bigg(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(D_{ij}\hat{e}_{j})\bigg)w_{k}\Psi_{k\mu}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}$ (119)
$(2):\hat{e}^{*}_{i}\bigg(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(D_{ij}\hat{e}_{j})\bigg)\frac{1}{2}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}$ (120)
$(3):\mathrm{i}\hat{e}^{*}_{i}\bigg(\partial_{\mu}\frac{\partial}{\partial K_{\nu}}(D_{ij}\hat{e}_{j})\bigg)w_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}$ (121)
@@
$\begin{split}\mathrm{i}\hat{e}^{*}_{i}\bigg(&\frac{\partial}{\partial K_{\nu}}D_{im}\Psi_{\nu k}w_{k}\frac{\partial}{\partial K_{l}}\hat{e}_{m}\partial_{l}(\mathcal{P})^{(0)}+\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}D_{im}\Psi_{\mu k}w_{k}\hat{e}_{m}\partial_{\nu}(\mathcal{P})^{(0)}\\ &+\frac{\partial}{\partial K_{\nu}}D_{im}\Psi_{lk}w_{k}\frac{\partial}{\partial K_{l}}\hat{e}_{m}\partial_{\nu}(\mathcal{P})^{(0)}\bigg)A^{(0)}.\end{split}$ (122)
@@
$\frac{1}{2}\hat{e}^{*}_{i}\left(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}D_{im}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}\hat{e}_{m}+2\frac{\partial}{\partial K_{\nu}}D_{im}\partial_{l}\partial_{\nu}(\mathcal{P})^{(0)}\frac{\partial}{\partial K_{l}}\hat{e}_{m}\right)A^{(0)}.$ (123)
@@
$\mathrm{i}\hat{e}^{*}_{i}\left(\frac{\partial}{\partial K_{\nu}}D_{im}w_{k}\partial_{k}\hat{e}_{m}\partial_{\nu}(\mathcal{P})^{(0)}+\partial_{k}D_{im}w_{k}\frac{\partial}{\partial K_{l}}\hat{e}_{m}\partial_{l}(\mathcal{P})^{(0)}\right)A^{(0)}.$ (124)
Since $\partial_{\mu}\frac{\partial}{\partial K_{\nu}}D_{ij}=0$, relabelling dummy indices gets us the expression in eq. 121. Thus, all the terms involving derivatives of the profile function add up to the following expression:
$\begin{split}\mathrm{i}&\Big(\partial_{\tau}(\mathcal{P})+w_{y}\frac{\mathrm{d}\hat{\bm{x}}}{\mathrm{d}\tau}\cdot\hat{\bm{y}}\partial_{w_{x}}(\mathcal{P})+w_{x}\frac{\mathrm{d}\hat{\bm{y}}}{\mathrm{d}\tau}\cdot\hat{\bm{x}}\partial_{w_{y}}(\mathcal{P})\Big)A^{(0)}+\mathrm{i}\hat{e}^{*}_{i}\left(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(D_{ij}\hat{e}_{j})\right)w_{k}\Psi_{k\mu}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}\\ &+\hat{e}^{*}_{i}\left(\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}(D_{ij}\hat{e}_{j})\right)\frac{1}{2}\partial_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}+\mathrm{i}\hat{e}^{*}_{i}\left(\partial_{\mu}\frac{\partial}{\partial K_{\nu}}(D_{ij}\hat{e}_{j})\right)w_{\mu}\partial_{\nu}(\mathcal{P})^{(0)}A^{(0)}.\end{split}$ (125)
@@
$\frac{\mathrm{d}}{\mathrm{d}\tau}(T_{\mu_{0}\mu_{1}\mu_{2}...\mu_{n}}):=\frac{\partial T_{\mu_{0}\mu_{1}\mu_{2}...\mu_{n}}}{\partial\tau}+\sum_{i=0}^{n}\sum_{j=0}^{2}\hat{x}_{j}^{\mu_{n}}\frac{\mathrm{d}\hat{x}_{j}^{\rho}}{\mathrm{d}\tau}T_{\mu_{0}...\mu_{n-1}\rho\mu_{n+1}...\mu_{n}}.$ (126)